xxcoffee:

what is 3(x +12) = 2 (2x +14) Work has to be shown

itsmehjay:

we got distribute the numbers outside of the parenthesis first :o

hhanan:

3x+36=4x+28 36=4x+28−3x 36=x+28 36−28=x 8=x x=8

itsmehjay:

i-

Yahaira:

Can someone help me?

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yahaira Can someone help me? \(\color{#0cbb34}{\text{End of Quote}}\) make a question then...

Yahaira:

​Together, teammates Pedro and Ricky got 2682 base hits last season. Pedro had 276 more hits than Ricky. How many hits did each player​ have?

Reece:

Hi, can someone please help me answer this question? Determine the exact value of each expression: cos (sin inverse (-square root of 3/2))

XioGonz:

Alright 3*(x+12)-(2*(2*x+14))=0 Pull out like factors: 2x + 14 = 2 • (x + 7) (3 • (x + 12)) - 4 • (x + 7) = 0 8 - x = 0 -x+8 = 0 Subtract 8 from both sides, which gives you: -x = -8 Multiply both sides of the equation by (-1) Hence, x = 8

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Alright 3*(x+12)-(2*(2*x+14))=0 Pull out like factors: 2x + 14 = 2 • (x + 7) (3 • (x + 12)) - 4 • (x + 7) = 0 8 - x = 0 -x+8 = 0 Subtract 8 from both sides, which gives you: -x = -8 Multiply both sides of the equation by (-1) Hence, x = 8 \(\color{#0cbb34}{\text{End of Quote}}\) thank you

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yahaira ​Together, teammates Pedro and Ricky got 2682 base hits last season. Pedro had 276 more hits than Ricky. How many hits did each player​ have? \(\color{#0cbb34}{\text{End of Quote}}\) Make ur own question by asking one on the upper left side of your screen. There post your question-

hhanan:

Make your own post and don't comment if you're not going to help:) @Reece @Yahaira

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxcoffee \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Alright 3*(x+12)-(2*(2*x+14))=0 Pull out like factors: 2x + 14 = 2 • (x + 7) (3 • (x + 12)) - 4 • (x + 7) = 0 8 - x = 0 -x+8 = 0 Subtract 8 from both sides, which gives you: -x = -8 Multiply both sides of the equation by (-1) Hence, x = 8 \(\color{#0cbb34}{\text{End of Quote}}\) thank you \(\color{#0cbb34}{\text{End of Quote}}\) Np (:

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan Make your own post and don't comment if you're not going to help:) @Reece @Yahaira \(\color{#0cbb34}{\text{End of Quote}}\) fr like im trying to keep from gettig a warning or ban

Yahaira:

Sorry, this is my first time using this.

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yahaira Sorry, this is my first time using this. \(\color{#0cbb34}{\text{End of Quote}}\) make ur own question ur goiing to make me get a warnin that i dont need rn

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yahaira Sorry, this is my first time using this. \(\color{#0cbb34}{\text{End of Quote}}\) It's alright Welcome to QC Ask your Question here

1 attachment
Yahaira:

ok